OpenStudy (anonymous):

if f = 158... c=(5/9)(F-32) answer for c

Directrix (directrix):

c = (5/9) ( F-32 ) C = (5/9) ( (158) - 32) ) C = 70 Questions?

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

it says its wrong when i put 70 in...

OpenStudy (anonymous):

it says f eaquals 158F when 70C (thats the first part) then it says solve for C from the equation i posted..

OpenStudy (anonymous):

can i send it to you word for word?

Directrix (directrix):

The F = 158 goes to C = 70. Check out this online calculator. http://www.wbuf.noaa.gov/tempfc.htm

OpenStudy (anonymous):

the formula C= 5over9(F-32) is used to find temp, C, in degrees Celsius for a given temp expressed in Fahrenheit, F. solve the formual for f. use this equation to find the equavalent F temp for 70 degrees C. the formula is F=____, and 70 degrees C =___ degrees F

OpenStudy (anonymous):

thats word for word

Directrix (directrix):

Okay. We'll do that and then check.

OpenStudy (anonymous):

i put 158 in the F blank and its telling me 70 would be the wrong answer for the C blank

Directrix (directrix):

Don't put anything else in there until we finish the problem.

Directrix (directrix):

Solve the formual for F: C = (5/9) ( F-32 ) (9/5) C = (9/5) ( (5/9) ( F - 32)

Directrix (directrix):

(9/5) C = (9/5) ( (5/9) ( F - 32) (9/5) C = ( F - 32 ) (9/5) C + 32 = F F = (9/5) C + 32

Directrix (directrix):

I have to check that - don't enter anything. We have two more calculations.

Directrix (directrix):

Now, the second part.

OpenStudy (anonymous):

F= 158...

Directrix (directrix):

The question says to put in 70 for C.

OpenStudy (anonymous):

the first part is finiding F then the second part is "and 70 degrees Celcius = blank degrees F

Directrix (directrix):

F = (9/5) (70) + 32 F = 158

Directrix (directrix):

70 degrees Celsius = 158 degrees Farenheit

Directrix (directrix):

Is there a third part?

OpenStudy (anonymous):

no that was the ? word for word

OpenStudy (anonymous):

oh i see i was suppose to put the equation in for F blank then 158 for the second part thank you